OpenStudy (anonymous):

Fan and medal whoever answers Ms. Malek is planning a breakfast meeting for 45 people. She wants at least one donut per person, but donuts are sold in boxes of a dozen. What is the least number of boxes she should buy?

OpenStudy (anonymous):

48

OpenStudy (anonymous):

if he buys 3 dozens (3*12=36donuts) he won't have enough donuts, but if he buys 4 (4*12=48) he'll have 48, which is 3 more, but enough all the people.

OpenStudy (anonymous):

One dozen is twelve, so the lowest multiple of 12 that is at least 45 is 48; hence, 12 * 4, so four boxes. \[45 \le 12x \rightarrow x \ge 3.75 \rightarrow x = 4\]

OpenStudy (anonymous):

12x4 48 r3